In order for a chemical reaction to take place, the fahrenheit temperature of the reagents must be atleast 149.25 degrees Fahrenheit. Find the Celsius temperatures at which the reaction may occur. (F=9 over 5 C + 32)
Answer by rfer(16322)

C=5/9(f-32)
C=5/9(149.25-32)
C=5/9*117.25
C=586.25/9
C=65.14 degrees